OpenStudy (anonymous):

Differentiate the following function. y=[(x-2)(9x+4)]^3

OpenStudy (anonymous):

Here you use the chain rule. Have you learned that yet?

OpenStudy (anonymous):

d dxf(g(x)) = f 0 (g(x))g 0 (x)

OpenStudy (anonymous):

not yet, just the product and quotient rule

OpenStudy (anonymous):

use that rule dood

OpenStudy (anonymous):

compute whats in the brackets and then simplify until you have a regular polynomial

OpenStudy (anonymous):

\[6(x-2)^{2}(9x-7)(9x+4)^{2}\]

OpenStudy (anonymous):

is that the differentiation of the function?

OpenStudy (anonymous):

simplify as much as possible

OpenStudy (anonymous):

Use the chain rule! It will take way too long using the product rule.

OpenStudy (anonymous):

I haven't learned the chain rule yet.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y=\left[(x-2)(9x+4) \right]^3 }\) when you take the derivative (dy/dx) (of function y, with respect to x) on both sides, the left side is just dy/dx denoting that it (what is on the right side) is the 1st derivative of the function. And the right side is as follows.... \(\large\color{black}{ \displaystyle \frac{ dy}{dx}=3\times \left[(x-2)(9x+4) \color{white}{\int}\right]^{3-1} \times \left[ \frac{ d}{dx} \left(~(x-2)(9x+4)~\right)\right]}\) so we are applying the power rule to the entire argument, and if the derivative of the inner argument is not 1, then we apply the chain rule.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{ dy}{dx}=3\times \left[(x-2)(9x+4) \color{white}{\int}\right]^{3-1} \times \left[ \frac{ d}{dx} \left(~(x-2)(9x+4)~\right)\right]}\) what will this be? ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ Well: \(\large\color{black}{ \displaystyle \frac{ d}{dx} \left(~(x-2)(9x+4)~\right)}\) product rule, (we get ) \(\large\color{black}{ \displaystyle (9x+4)\frac{ d}{dx} (x-2)~+~(x-2)\frac{ d}{dx} (9x+4)}\) \(\large\color{black}{ \displaystyle (9x+4)1~+~(x-2)9}\) \(\large\color{black}{ \displaystyle 9x+4~+~9x-18}\) \(\large\color{black}{ \displaystyle 18x-14}\) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ So, \(\large\color{black}{ \displaystyle \frac{ dy}{dx}=3\times \left[(x-2)(9x+4) \color{white}{\int}\right]^{3-1} \times \left[ \frac{ d}{dx} \left(~(x-2)(9x+4)~\right)\right]}\) \(\large\color{black}{ \displaystyle \frac{ dy}{dx}=3\times \left[(x-2)(9x+4) \color{white}{\int}\right]^{3-1} \times \left[ 18x-14\right]}\) \(\large\color{blue}{ \displaystyle \frac{ dy}{dx}=3\times \left[(x-2)(9x+4) \color{white}{\int}\right]^{2} \times \left[ 18x-14\right]}\) this in blue is the derivative

OpenStudy (noelgreco):

It will be easier if you expand the product inside the brackets to yield quadratic, the follow the exponent and chain rules.

OpenStudy (solomonzelman):

you can use logarithmic differentiation as well:)

OpenStudy (solomonzelman):

it will be quite well coming out